OpenStudy (anonymous):

If the circumference of a circle is 50pi , calculate the area of the circle.

OpenStudy (angela210793):

A=pi*r^2

OpenStudy (anonymous):

i think that its 25pi but im not sure

OpenStudy (lalaly):

circumference = 2pi*r =50pi 2r=50 r=25 area=pi*r^2 = pi*25^2

OpenStudy (angela210793):

P=50pi=2pir--->r=25 A=25^2pi=625pi

OpenStudy (ksaimouli):

50p/2=r

OpenStudy (ksaimouli):

50p/2r2

OpenStudy (ksaimouli):

25p/2

OpenStudy (ksaimouli):

p*252

OpenStudy (anonymous):

the forumla for the area of the circle is \[C=2\pi r\] rearrange for the radius, \[r = C/2\pi\], plug the values and you get radius 25. now the area of the circle can be calulcated as \[A=\pi r^2 = \pi 25^2 = 1963.49 m^{2}\]